# Electric birefringence in Euler-Heisenberg pseudo-electrodynamics

### v1

@@ I Introduction
Field theories in $1+2$ dimensions have attracted great interest due to it application in condensed matter physics. Materials like graphene in connection with the quantum Hall effect [1, 2, 3, 4, 5, 6, 7], topological isolators [8, 9, 10, 11, 12], superconductivity in layered materials [13, 14, 15, 16, 17] have showed that quantum field theories (QFTs) provide an excellent theoretical description for experimental results in material physics. In Dirac materials, electrons move in Fermi velocity through a semi-relativistic dynamics are good material class to test QFTs [18]. The interactions between electrons in planar materials are mediated by a gauge field theory known as pseudo-electrodynamics (PED) [19]. The PED is a non-local Abelian gauge theory obtained confining the classical sources of the usual Maxwell electrodynamics to a spatial plane, that results in the dimensional reduction to $1+2$ dimensions. When the PED is coupled to fermions through a $U(1)$ local gauge invariance, the model is called pseudo-quantum electrodynamics (PQED). Important properties of the PED were demonstrated as the canonical quantization, causality and unitarity [20, 21]. Several extensions of the PED and PQED also were studied through the QFT approach, as the addition of Chern-Simons topological term [22, 23], the Proca pseudo-electrodynamics [24, 25, 26, 27], the Lee-Wick pseudo-electrodynamics [28, 29, 30]. In all these references different approach are discussed to include mass to the pseudo-gauge field.
In parallel to the success of the PQED, investigations of non-linear effects from the quantum electrodynamics applied to Dirac material emerges from polarized vacuum Dirac sea [31, 32]. Non-linear electrodynamics (NLEDs) emerged for the first time with M. Born and L. Infeld to explain the static electric field of a like-point charge at origin [33]. Some years after, H. Euler and W. Heisenberg obtained the fermionic effective action under an external EM field at one loop in the QED [34]. In present days, others NLEDs has been proposed with several applications in many research areas [35, 36, 37, 38, 39, 40, 41]. The observation of non-linear effects also appear in the Hall effect [42, 43, 44], and in topological photonics [45]. The results of these references motivate us to explore the non-linear electrodynamics that arises from the radiative corrections at one loop in the PQED.
@@
We consider the natural units with $\hbar=c=1$ throughout paper. The metric signature is $\eta_{\bar{\mu}\bar{\nu}}=\mbox{diag}(+1,-1,-1)$, in which the bar index $\bar{\mu}=\{0,1,2\}$ are adopted for vectors and tensors in $1+2$ dimensions. The electric and magnetic fields have the following conversions of $\mbox{eV}^{2}$ to V/cm and Tesla (T) in natural units : $1\,\mbox{eV}^{2}=4.4\times 10^{3}\,\mbox{V/cm}$ and $1\,\mbox{eV}^{2}=1.47\times 10^{-3}\,\mbox{T}$, respectively.
@@ II The Euler-Heisenberg pseudo-electrodynamics
where $\psi$ is a Dirac spinor of two components, $\Delta=m\,v_{F}^{2}$ is the gap over two of the planar material with $m=0.5$ MeV being the electron mass, $\Gamma^{\bar{\mu}}=(\gamma^{0},v_{F}\gamma^{i})\,(i=1,2)$ are the Dirac matrices in which the Fermi velocity $(v_{F})$ emerges in the spatial component of the $\gamma^{\bar{\mu}}$-matrices, that satisfy the relation $\gamma^{\bar{\mu}}\gamma^{\bar{\nu}}=\eta^{\bar{\mu}\bar{\nu}}+i\,\sigma^{\bar{\mu}\bar{\nu}}$, and $\bar{\psi}=\psi^{\dagger}\,\gamma^{0}$ is the adjunct field. The covariant derivative operator is $D_{\bar{\mu}}=\partial_{\bar{\mu}}+ie\,A_{\bar{\mu}}$, in which the fundamental charge $(e)$ is related to fine structure constant $(\alpha=1/137)$ by $e^{2}=4\pi\,\alpha\,v_{F}$ in natural units. The notation for derivative operator is such that $\partial_{\bar{\mu}}=(\partial_{t},\partial_{x},\partial_{y})$, and the D’Alambertian operator in $1+2$ dimensions is $\bar{\Box}=\partial_{\bar{\mu}}\partial^{\bar{\mu}}=\partial_{t}^{2}-\partial_{x}^{2}-\partial_{y}^{2}$. The matrices $\gamma^{\bar{\mu}}=(\gamma^{0},\gamma^{1},\gamma^{2})$ are represented in terms of the Pauli matrices $(\sigma_{1},\sigma_{2},\sigma_{3})$ as
@@
We derivative (14) in relation to $\Delta^{2}$ to obtain the expression
@@
that substituting in the integral (27), the first term is divergent in $s=0$. Therefore, we need to introduce a renormalization scheme in the effective lagrangian (27). The regulator parameter $(s_{0})$ is introduced in the lower limit of the $s$-integral, and the counter-term needed to renormalization is given by
@@
in which $f(v_{F})=1-6v_{F}^{2}$ and $g(v_{F})=v_{F}^{4}-v_{F}^{2}+3/20$ are functions of the Fermi velocity. As example, for the graphene, the Fermi velocity is $v_{F}=0.0033$, such that these functions can be approximated by $f\simeq 0.99$ and $g\simeq 0.15$. The effective lagrangian is
@@
$E_{c}=\frac{\Delta^{2}}{e\,v_{F}}=5.16\times 10^{5}\,\mbox{eV}^{2}=2.27\times 10^{9}\,\mbox{V}/\mbox{cm}\;,$ (35)
@@
$B_{c}=\frac{E_{c}}{v_{F}}=1.56\times 10^{8}\,\mbox{eV}=2.29\times 10^{5}\,\mbox{T}\;,$ (36)
where $v_{F}=0.003$, $e=\sqrt{4\pi\alpha v_{F}}=0.017$ and $\Delta=5.44$ eV for the example of the graphene, in which the effects of the EM field are sensible up to this scale. Therefore, non-linear effects ruled by the quartic term in the tensor $G^{\bar{\mu}\bar{\nu}}$ of the effective lagrangian shows the emergence of a non-linear electrodynamics in $1+2$ dimensions. The PEHED (34) remains gauge invariant, but the presence of the Fermi velocity in the non-linear term breaks the Lorentz symmetry. We will investigate some of the effects caused by the non-linearity in the next section.
@@ III Dispersion relations of the PEH in a external EM field
in which are real quantities in the frequency range of $\omega<|{\bf k}|$. The solution (47a) is the usual vacuum electric permittivity, while that the results (47b) and (48) show that the material medium is dispersive due to dependence with the wave vector $(\bf k)$ and $\omega$-frequency. In fact, this dependence is exclusively due to non-locality of the EHPED, and also to the background fields. Turning off the background fields, the results reduce to the usual vacuum, i.e., $\varepsilon_{1}=\varepsilon_{2}=\mu=1$. The non-linear effects of (47b) and (48) are relevant for EM background of
$|{\bf E}_{0}|\sim\sqrt{\frac{64\pi\Delta^{5}}{e^{4}\,g(v_{F})\,\omega}}\simeq 1.57\times 10^{10}\,\mbox{V}/\mbox{cm}\,\left[\frac{5.44\,\mbox{eV}}{\omega}\right]^{1/2}\!\!,$
$v_{F}^{2}B_{0}\sim\sqrt{\frac{64\pi\Delta^{5}}{e^{4}\,g(v_{F})\,\omega}}\simeq 5.26\times 10^{3}\,\mbox{T}\,\left[\frac{5.44\,\mbox{eV}}{\omega}\right]^{1/2}\!\!,\hskip 25.6073pt$ (49b)
when $v_{F}\ll 1$, and we have used the conversion of $\mbox{MeV}^{2}$ to V/cm and Tesla (T), respectively.
Substituting the Faraday-Lenz in the Ampère-Maxwell law in (43), the wave equation for $e_{0i}$ is $M_{ij}e_{0j}=0$, where the symmetric wave matrix $M_{ij}$ is
@@ IV Birefringence in PEHED
In this section, we discuss the birefringence phenomenon in the case of the pseudo-EH ED. Since the EM plane wave propagates over the plane $\cal XY$}, the most general wave polarization is ${\bf e}_{0}=(e_{0x}\,,\,e_{0y})$, as well as, the wave vector is ${\bf k}=(k_{x}\,,\,k_{y})$. The birefringence analysis starts with particular considerations on the these quantities in which the propagation direction is ${\bf k}=(k,0)$, and the wave polarization is linear on the $\cal X$}-direction, or on the $\cal Y$}-direction. If ${\bf e}_{0}=(e_{0x}\,,\,0)$, the wave equation $M_{11}\,e_{0x}=0$ reproduces a parallel refractive index $n_{\parallel}$ as solution, and when we have the situation of ${\bf e}_{0}=(0\,,\,e_{0y})$, the wave equation $M_{22}\,e_{0y}=0$ yields the solution of $n_{\perp}$. The birefringence is defined by
@@
In the case of a pure magnetic background field $({\bf E}_{0}={\bf 0})$, the wave matrix $M_{ij}$ shows that the birefringence is null in $1+2$ dimensions, i.e., $\Delta n_{B}=0$.
When $B_{0}=0$, we consider the situation of the electric background field on the $\cal Y$}-direction, with $^{\bf E}_{0}=E_{0}\,\hat{{\bf y}}$, and the wave equation $M_{11}\,e_{0x}=0$ yields the parallel refractive index $n_{\parallel}=1$ with the wave propagation direction. If the wave polarization is on $\cal Y$}, the wave equation $M_{22}\,e_{0y}=0$ has the solution
$n_{\perp}=\sqrt{1+\frac{e^{8}g^{2}(v_{F})\,\omega^{2}E_{0}^{4}}{4096\pi^{2}\Delta^{10}}}\simeq 1+\frac{e^{8}g^{2}(v_{F})\,\omega^{2}E_{0}^{4}}{8192\pi^{2}\Delta^{10}}\;.$ (65)
@@
The result shows that the birefringence is proportional to wave frequency to square, and also is proportional to the electric background to fourth power. If we substitute the values of $v_{F}=0.003$, $e=\sqrt{4\pi\alpha v_{F}}=0.017$ and $\Delta=5.44$ eV, we obtain the result for any electric background :
$\frac{\Delta n_{E}}{\omega^{2}E_{0}^{4}}\simeq 1.019\times 10^{-28}\,\mbox{eV}^{-10}\;.$ (67)
Using the values of $E_{0}=3.58\times 10^{6}\,\mbox{eV}^{2}$ (from eq. (49) ) and $\omega\sim 5.44\,\mbox{eV}$, we obtain the result
@@ V Conclusions
In this paper, the effective action of the fermion sector in the pseudo-quantum electrodynamics (PQED) is calculated to generate the so called Euler-Heisenberg pseudo-electrodynamics (PEHED). It is a non-linear electrodynamics defined on the planar space of $1+2$ dimensions. The Chern-Simons (CS) topological term is added to achieve a complete PED gauge invariant model. Since the fermions sector has a Lorentz symmetry violation in the PQED, this breaking Lorentz symmetry is transferred to the effective non-linear lagrangian. As consequence, the PEHED breaks the Lorentz symmetry due to presence of the Fermi velocity, that depends on the material nature. We explore the effects of the PEHED when the model is submitted to an external electromagnetic background, in which the fields are considered uniform and constant. The electric permittivity tensor and magnetic permeability are calculated in terms of the EM external field. These results show that the planar material is a dispersive medium in which one of the eigenvalues of the electric permittivity tensor is function of the wave vector $({\bf k})$, and of the frequency $(\omega)$. The magnetic permeability is a scalar in $1+2$ dimensions, that also shows the dependence with $\bf k$} and $\omega$. Posteriorly, we obtain the solutions for the refractive index of this planar material medium in the presence of a magnetic background field, and also for the case of only an electric background field. The results show three solutions for the refractive index : The first is inversely proportional to wave frequency times the squared magnetic background field; the second and third solutions provide complex solutions, where the imaginary parts are interpreted as wave absorptions that depend on the CS parameter and on the radiative corrections emerged from the fermionic effective action. The case of a pure electric background field $({\bf E}_{0})$ is similar to magnetic case, but in the first solution, the refractive index depends on the angle that ${\bf E}_{0}$ does with $\bf k$}. When ${\bf E}_{0}$ is parallel to $\bf k$}-direction, the solutions are reduced to the refractive index of the CS pseudo-electrodynamics. Also in the electric case, two solutions exhibit real and imaginary parts with the wave absorption that are non null if ${\bf E}_{0}$ is not parallel to $\bf k$}-direction. These results for the refractive index are compatible with the complex refractive index in the graphene, see the ref. [49].
@@
For end, the birefringence is studied with the help of the wave equation in the EM background field. When only the external magnetic field acts on the planar medium, there is no birefringence. This phenomenon appears only when we consider the electric background field. The non-local structure of the PEHED leads to a birefringence proportional to $\omega^{2}E_{0}^{4}$, whose effect is of the order of $\sim 1.019\times 10^{-28}\,\mbox{eV}^{-10}$. If we use a typical electric field of $E_{0}=10^{6}\,\mbox{eV}^{2}$ in a frequency of $\omega\sim 5\,\mbox{eV}$, the permittivity and permeability of the planar medium is relevant, and the electric birefringence is estimated by $\Delta n_{E}\simeq 0.49$. In the context of the dimensional reduction that led to PED and of the optical properties, others non-linear electrodynamics known in the literature under the prescription of the EM external fields can be investigated in a forthcoming project.

### v2

@@ I Introduction
Field theories in $1+2$ dimensions have attracted great interest due to it application in condensed matter physics. Materials like graphene in connection with the quantum Hall effect [1, 2, 3, 4, 5, 6, 7], topological isolators [8, 9, 10, 11, 12], superconductivity in layered materials [13, 14, 15, 16, 17] have showed that quantum field theories (QFTs) provide an excellent theoretical description for experimental results in material physics. In Dirac materials, electrons move at Fermi velocity through a semi-relativistic dynamics are good material class to test QFTs [18]. The interactions between electrons in planar materials are mediated by a gauge field theory known as pseudo-electrodynamics (PED) [19]. The PED is a non-local Abelian gauge theory obtained confining the classical sources of the usual Maxwell electrodynamics to a spatial plane, that results in the dimensional reduction to $1+2$ dimensions. When the PED is coupled to fermions through a $U(1)$ local gauge invariance, the model is called pseudo-quantum electrodynamics (PQED). Important properties of the PED were demonstrated as the canonical quantization, causality and unitarity [20, 21]. Several extensions of the PED and PQED also were studied through the QFT approach, as the addition of Chern-Simons topological term [22, 23], the Proca pseudo-electrodynamics [24, 25, 26, 27], the Lee-Wick pseudo-electrodynamics [28, 29, 30], and supersymmetry [31]. In all these references different approaches are discussed to include mass to the pseudo-gauge field.
In parallel to the success of the PQED, investigations of non-linear effects from the quantum electrodynamics applied to Dirac material emerges from polarized vacuum Dirac sea [32, 33]. Non-linear electrodynamics (NLEDs) emerged for the first time with M. Born and L. Infeld to explain the static electric field of a point charge at origin [34]. Some years after, H. Euler and W. Heisenberg obtained the fermionic effective action under an external EM field at one loop in the QED [35]. In present days, others NLEDs has been proposed with several applications in many research areas as Cosmology, alternative Gravitation theories, material physics and others, see the references [36, 37, 38, 39, 40, 41, 42, 43, 44]. The observation of non-linear effects also appear in the Hall effect [45, 46, 47], and in topological photonics [48]. Effects of QED has been studied in graphene, as the Schwinger effect [49, 50, 51, 52, 53, 54, 55]. The results of all these references motivate us to explore the non-linear ED that arises from the radiative corrections at one loop in the PQED. This non-linear and planar ED has not been explored in the literature yet.
@@
We consider the natural units with $\hbar=c=1$ throughout paper. The metric signature is $\eta_{\bar{\mu}\bar{\nu}}=\mbox{diag}(+1,-1,-1)$, in which the bar index $\bar{\mu}=\{0,1,2\}$ are adopted for vectors and tensors in $1+2$ dimensions. The electric and magnetic fields have the following conversions of $\mbox{eV}^{2}$ to V/cm and Tesla (T) in natural units : $1\,\mbox{eV}^{2}=1.54\times 10^{4}\,\mbox{V/cm}$ and $1\,\mbox{eV}^{2}=692.5\,\mbox{T}$, respectively.
@@ II The Euler-Heisenberg pseudo-electrodynamics
where $\psi$ is a Dirac spinor of two components, $\Delta=0.1$ eV is the gap over two of the planar material. For many examples of Dirac isolators and graphene, we have the gap of $2\Delta\simeq 2$ eV. $\Gamma^{\bar{\mu}}=(\gamma^{0},v_{F}\gamma^{i})\,(i=1,2)$ are the Dirac matrices in which the Fermi velocity $(v_{F})$ emerges in the spatial component of the $\gamma^{\bar{\mu}}$-matrices, that satisfy the relation $\gamma^{\bar{\mu}}\gamma^{\bar{\nu}}=\eta^{\bar{\mu}\bar{\nu}}+i\,\sigma^{\bar{\mu}\bar{\nu}}$, and $\bar{\psi}=\psi^{\dagger}\,\gamma^{0}$ is the adjunct field. The covariant derivative operator is $D_{\bar{\mu}}=\partial_{\bar{\mu}}+ie\,A_{\bar{\mu}}$, in which the fundamental charge $(e)$ is related to fine structure constant $(\alpha=1/137)$ by $e^{2}=4\pi\,\alpha\,v_{F}$ in natural units. The notation for derivative operator is such that $\partial_{\bar{\mu}}=(\partial_{t},\partial_{x},\partial_{y})$, and the D’Alambertian operator in $1+2$ dimensions is $\bar{\Box}=\partial_{\bar{\mu}}\partial^{\bar{\mu}}=\partial_{t}^{2}-\partial_{x}^{2}-\partial_{y}^{2}$. The matrices $\gamma^{\bar{\mu}}=(\gamma^{0},\gamma^{1},\gamma^{2})$ are represented in terms of the Pauli matrices $(\sigma_{1},\sigma_{2},\sigma_{3})$ as
@@
We calculate the derivative of (14) in relation to $\Delta^{2}$ to obtain the expression
@@
that substituting in the integral (27), the first term is divergent in $s=0$. Therefore, we need to introduce a renormalization scheme in the effective lagrangian (27). The regulator parameter $(s_{0})$ is introduced in the lower limit of the $s$-integral, and the counter-term needed for renormalization is given by
@@
in which $f(v_{F})=1-6v_{F}^{2}$ and $g(v_{F})=v_{F}^{4}-v_{F}^{2}+3/20$ are functions of the Fermi velocity. As example, for the graphene, the Fermi velocity in natural units is $v_{F}=0.0033$, such that these functions can be approximated by $f\simeq 0.99$ and $g\simeq 0.15$. The effective lagrangian is
@@
$E_{c}=\frac{\Delta^{2}}{e\,v_{F}}=174.17\,\mbox{eV}^{2}=2.68\times 10^{6}\,\mbox{V}/\mbox{cm}\;,$ (35)
@@
$B_{c}=\frac{E_{c}}{v_{F}}=5.2\times 10^{4}\,\mbox{eV}^{2}=3.6\times 10^{7}\,\mbox{T}\;,$ (36)
where $v_{F}=0.003$, $e=\sqrt{4\pi\alpha v_{F}}=0.017$ and $\Delta=0.1$ eV for the example of the graphene, in which the effects of the EM field are sensible up to this scale. The electric critical field is in same order of the typical values of the electric fields in Dirac materials [32]. For the regime of validity of the (34), the electric and magnetic fields must satisfy the conditions $|{\bf E}|\ll E_{c}$ and $|{\bf B}|\ll B_{c}$, respectively.
Therefore, non-linear effects ruled by the quartic term in the tensor $G^{\bar{\mu}\bar{\nu}}$ of the effective lagrangian shows the emergence of a non-linear electrodynamics in $1+2$ dimensions. The PEHED (34) remains gauge invariant, but the presence of the Fermi velocity in the non-linear term breaks the Lorentz symmetry. We will investigate some of the effects caused by the non-linearity in the next section.
@@ III Dispersion relations of the PEH in a external EM field
in which are real quantities in the frequency range of $\omega<|{\bf k}|$. The solution (47a) is the usual vacuum electric permittivity, while that the results (47b) and (48) show that the material medium is dispersive due to presence of the wave vector $(\bf k)$ and $\omega$-frequency. In fact, this dependency is exclusively due to non-locality of the EHPED, and also to the background fields. Turning off the background fields, the results reduce to the usual vacuum, i.e., $\varepsilon_{1}=\varepsilon_{2}=\mu=1$. The non-linear effects of (47b) and (48) are relevant for EM background of
$|{\bf E}_{0}|\sim\sqrt{\frac{64\pi\Delta^{5}}{e^{4}\,g(v_{F})\,\omega}}\simeq 1.86\times 10^{7}\,\mbox{V}/\mbox{cm}\,\left[\frac{0.1\,\mbox{eV}}{\omega}\right]^{1/2}\!\!,$
$v_{F}^{2}B_{0}\sim\sqrt{\frac{64\pi\Delta^{5}}{e^{4}\,g(v_{F})\,\omega}}\simeq 8.3\times 10^{5}\,\mbox{T}\,\left[\frac{0.1\,\mbox{eV}}{\omega}\right]^{1/2}\!\!,\hskip 25.6073pt$ (49b)
when $v_{F}\ll 1$, and we have used the conversion of $\mbox{MeV}^{2}$ to V/cm and Tesla (T), respectively. For values of EM background compatible with the critical fields (35) and (36), we impose the lower bound $\omega>0.1$ eV on the wave frequency.
Substituting the Faraday-Lenz in the Ampère-Maxwell law in (43), the wave equation $M_{ij}e_{0j}=0$ for a general wave polarization $e_{0j}$, in which the symmetric wave matrix $M_{ij}$ is
@@ IV Birefringence in PEHED
In this section, we discuss the birefringence phenomenon in the case of the pseudo-EH ED. Since the EM plane wave propagates over the plane $\cal XY$}, the most general wave polarization is ${\bf e}_{0}=(e_{0x}\,,\,e_{0y})$, as well as, the wave vector is ${\bf k}=(k_{x}\,,\,k_{y})$. The birefringence analysis starts with particular considerations on the direction of the external field (magnetic or electric) and the wave polarization on the $\cal X$}-direction, or on the $\cal Y$}-direction. Since the PEHED is an planar theory defined on a space of $1+2$ dimensions, the external magnetic field is on the $\cal Z$}-direction and the wave polarization vectors are on the bidimensional plane. If we change the direction of the wave polarization vectors, these vectors are perpendiculars to the external magnetic field in any situation. Thereby, the birefringence associated with the magnetic background field does not appear in $1+2$ dimensions. The case with electric background field is interesting due to electric vector is defined on the bidimensional plane. When $B_{0}=0$, we consider the situation of the electric background field on the $\cal Y$}-direction, with $^{\bf E}_{0}=E_{0}\,\hat{{\bf y}}$. If we set the wave polarization as ${\bf e}_{0}=(e_{0x}\,,\,0)$, the wave equation is $M_{11}\,e_{0x}=0$, and it reproduces a perpendicular refractive index $n_{\perp}$ as solution. When we have the situation of a linear wave polarization on $\cal Y$}-direction, with ${\bf e}_{0}=(0\,,\,e_{0y})$, the wave equation is $M_{22}\,e_{0y}=0$, in which it yields the solution of the parallel refractive index $n_{\parallel}$. The birefringence is defined by the subtraction :
@@
The wave equation $M_{11}\,e_{0x}=0$ yields the perpendicular refractive index $n_{\perp}=1$. If the wave polarization is on $\cal Y$}-direction, the wave equation $M_{22}\,e_{0y}=0$ has the solution
$n_{\parallel}=\sqrt{1+\frac{e^{8}g^{2}(v_{F})\,\omega^{2}E_{0}^{4}}{4096\pi^{2}\Delta^{10}}}\simeq 1+\frac{e^{8}g^{2}(v_{F})\,\omega^{2}E_{0}^{4}}{8192\pi^{2}\Delta^{10}}\;.$ (65)
@@
The result shows that the birefringence is proportional to wave frequency to square, and also is proportional to the electric background to fourth power. If we substitute the values of $v_{F}=0.003$, $e=\sqrt{4\pi\alpha v_{F}}=0.017$ and $\Delta=0.1$ eV, we obtain the result for any electric background :
$\frac{\Delta n_{E}}{\omega^{2}E_{0}^{4}}\simeq 2.33\times 10^{-11}\,\mbox{eV}^{-10}\;.$ (67)
Using the values of $E_{0}=1.2\times 10^{3}\,\mbox{eV}^{2}$ (from eq. (49) ) and $\omega\sim 0.1\,\mbox{eV}$, we obtain the result
@@ V Conclusions
In this paper, the effective action of the fermion sector in the pseudo-quantum electrodynamics (PQED) is calculated to generate the so called Euler-Heisenberg pseudo-electrodynamics (PEHED). It is a non-linear electrodynamics defined on the planar space of $1+2$ dimensions. The Chern-Simons (CS) topological term is added to achieve a complete PED gauge invariant model. The fermions sector has a Lorentz symmetry violation in the PQED, this Lorentz symmetry breaking is transferred to the effective non-linear lagrangian. As consequence, the PEHED breaks the Lorentz symmetry due to presence of the Fermi velocity, that depends on the nature of the materia medium. We explore the effects of the PEHED when the model is submitted to an external electromagnetic background, where these fields are considered uniform and constant. The electric permittivity tensor and magnetic permeability are calculated in terms of the EM external field. These results show that the planar material is a dispersive medium in which one of the eigenvalues of the electric permittivity tensor is function of the wave vector $({\bf k})$, and of the frequency $(\omega)$. The magnetic permeability is a scalar in $1+2$ dimensions, that also is function of $\bf k$} and $\omega$. Posteriorly, we obtain the solutions for the refractive index of this planar material medium in the presence of a magnetic background field, and also for the case of only an electric background field, when the wave has a general polarization. The results show three solutions for the refractive index : The first is inversely proportional to wave frequency times the squared magnetic background field; the second and third solutions provide complex solutions, where the imaginary parts are interpreted as wave absorptions that depend on the CS parameter, and on the radiative corrections emerged from the fermionic effective action. The case of a pure electric background field $({\bf E}_{0})$ is similar to magnetic case, but in the first solution, the refractive index depends on the angle that ${\bf E}_{0}$ does with $\bf k$}. When ${\bf E}_{0}$ is parallel to $\bf k$}-direction, the solutions are reduced to the refractive index of the CS pseudo-electrodynamics. Also in the electric case, two solutions exhibit real and imaginary parts with the wave absorption that are non null if ${\bf E}_{0}$ is not parallel to $\bf k$}-direction. These results for the refractive index are compatible with the complex refractive index in the graphene, see the ref. [59].
@@
For end, the birefringence is studied with the help of the wave equation in the EM background field. Since the external magnetic field has only component perpendicular to the planar space, the magnetic birefringence does not appear in $1+2$ dimensions. This phenomenon appears only when we consider the electric background field, where we study the birefringence for linear wave polarization. The non-local structure of the PEHED leads to a birefringence proportional to $\omega^{2}E_{0}^{4}$, whose effect is of the order of $\sim 2.33\times 10^{-11}\,\mbox{eV}^{-10}$. If we use a typical electric field of $E_{0}=10^{3}\,\mbox{eV}^{2}$ in a frequency of $\omega\sim 0.1\,\mbox{eV}$, the permittivity and permeability of the planar medium is relevant, and the electric birefringence is estimated by $\Delta n_{E}\simeq 0.48$. In the context of the dimensional reduction that led to PED and of the optical properties, others non-linear electrodynamics known in the literature under the prescription of the EM external fields can be investigated in a forthcoming project.